OpenStudy (thejoker):

Given Line Segment CD as shown with Line Segment CA~=(That is a congruent symbol)Line Segment BD and CD=98. If CB=3(x+8) and AD=4(x+3), find AB

OpenStudy (thejoker):



OpenStudy (thewafflebro):

@TheSmartOne

OpenStudy (nikato):

let's call AB, y so you can set up an equation CB + AD - AB = CD (you subtract AB because it is counted twice )

OpenStudy (thejoker):

The answers could be 12, 22, 36, 10, 42, 18, 6, 8, 30, 2, or 124

OpenStudy (nikato):

so 3(x+8) +4(x+3) -y =98

OpenStudy (thejoker):

I tried that and kept getting 99 instead, but when I put 1 as the y It came to be 0 putting it back in

OpenStudy (nikato):

did you solve for x first?

OpenStudy (nikato):

since CA is congruent to BD, you can set this up 3(x+8) = 4(x+3) what did you get for x?

OpenStudy (thejoker):

x=5

OpenStudy (nikato):

maybe check again? i got 12

OpenStudy (thejoker):

that is what I meant I hit 5 instead

OpenStudy (nikato):

ok. cool now plug this into the equation i set in the beginning and solve for y

OpenStudy (thejoker):

y=22

OpenStudy (nikato):

yes. 22 is coreect

OpenStudy (thejoker):

thanks